# Serenade: A Singing Style Conversion Framework Based On Audio Infilling

### v1

@@ IV-B Model architecture
For the prior encoder and GST style encoder, we used the same model configurations as in NU-SVC [17]. For the UNet used in the flow-matching model, we base it on the 1D UNet architecture proposed in Matcha-TTS [23] for high-quality synthesis. However, to further improve style modeling, we also conditioned the style embeddings after every residual block and increased the attention head dimension to 256 and the channel size to 512.
@@ IV-C Evaluation tests and baseline system
For subjective evaluations, we recruited 106 listeners and conducted the 5-scale mean opinion score test evaluating naturalness (N-MOS). Following the SVCC 2023 protocol, the definition of naturalness was on the listeners’ discretion. We also conducted a 4-scale XAB singing style similarity test, by showing listeners the source (A) and target (B) phrases and rate which was more similar to the converted phrase (X). Choosing B results in a score of 4, while choosing A results in a score of 1. X, A, and B are parallel and thus contain the same lyrics but with just different styles. In cases where the listener has to evaluate the ground truth’s similarity, we instead use a pitch-shifted version (of either half an octave up or down) of the source and target waveforms to ensure that listeners would not be easily able to distinguish the samples to be completely similar.
For the baseline system, we use NU-SVC [17], a DDPM-based SVC system, but replace the DDPM decoder with the flow-matching decoder. NU-SVC is similar to Serenade, but does not use the audio infilling task, and directly predicts the entire target mel-spectrogram from the conditioning features. For ablation studies, we investigate the use of audio and score MIDI, the use of large-scale pretraining, the use of the cyclic training, and the use of the SiFiGAN post-processing.
@@ V-A Effectiveness of Serenade’s audio infilling task
We show that the proposed Serenade SiFiGAN framework had the highest overall N-MOS in the proposed systems. Despite the use of cyclic training to improve similarity scores, the N-MOS in Serenade was still very low due to the out-of-tune singing. However, Serenade SiFiGAN helped alleviate this by using the original F0 and correcting out-of-tune singing, resulting in a much higher naturalness score. We also observed that there was a slight degradation in naturalness in Serenade SiFiGAN compared to NU-SVC. This is a downside of the SiFiGAN postprocessing scheme, as simply replacing the F0 with the source’s would have likely introduced small mismatches in WORLD features.
@@ V-B Ablation studies
We first investigate compared the use of score MIDI labels provided in GTSinger (Ablation 1) with the use of the audio-extracted MIDI using the pretrained network [19] (Ablation 2). We found that using audio-extracted MIDI was better than using score MIDI, primarily because we opted to use audio-extracted MIDI during inference to keep the system practical for SSC. Nevertheless, we still found that using audio-extracted MIDI, and being more correlated to the input waveform, was more helpful in normalizing the pitch.
@@ V-C Correlations between objective and subjective tests
Finding a well-correlated objective metric can help future researchers validate ideas before running expensive subjective tests. With no previous works on SSC, we found that reference-based metric FAD [24] using MERT [25] embeddings was the most correlated to the similarity scores, with a 0.93 correlation score. Nevertheless, large-scale evaluations with a multiple variety of systems still need to be conducted to find the best objective metric.
@@ VI Conclusions
We proposed Serenade, a novel framework to accomplish the SSC task and resolved the three main issues we found in this new task, namely: singing style modeling, source style disentanglement, and melody retention. We showed a more generalized SSC framework and better modeling in breath features due to the audio infilling. Next, we used a cyclic training strategy to further disentangle the singing style of the source singer. Finally, we were able to alleviate out-of-tune singing by introducing a SiFiGAN vocoder post-processing method. We conclude that the Serenade framework showed better similarity scores in breathy and mixed singing styles than the baseline system but with a slight degradation in naturalness to do so.
@@ Acknowledgment
This work was partly supported by JST CREST under Grant Number JPMJCR19A3.

### v2

@@ IV-B Model architecture
For the prior encoder and GST style encoder, we used the same model configurations as in NU-SVC [yamamoto2023svcc]. For the UNet used in the flow-matching model, we base it on the 1D UNet architecture proposed in Matcha-TTS [mehta2024matcha] for high-quality synthesis. However, to further improve style modeling, we also conditioned the style embeddings after every residual block and increased the attention head dimension to 256 and the channel size to 512.
@@ IV-C Evaluation tests and baseline system
For subjective evaluations, we recruited 106 listeners and conducted the 5-scale mean opinion score test evaluating naturalness (N-MOS). Following the SVCC 2023 protocol, the definition of naturalness was on the listeners’ discretion. Also following the SVCC 2023 evaluation, we also conducted a 4-scale XAB singing style similarity test, by showing listeners the source (A) and target (B) phrases and rate which was more similar to the converted phrase (X). Choosing B results in a score of 4, while choosing A results in a score of 1. X, A, and B are parallel and thus contain the same lyrics but with just different styles. In cases where the listener has to evaluate the ground truth’s similarity, we instead use a pitch-shifted version (of either half an octave up or down) of the source and target waveforms to ensure that listeners would not be easily able to distinguish the samples to be completely similar.
For the baseline system, we use NU-SVC [yamamoto2023svcc], a DDPM-based SVC system, but replace the DDPM decoder with the flow-matching decoder. NU-SVC is similar to Serenade, but does not use the audio infilling task, and directly predicts the entire target mel-spectrogram from the conditioning features. For ablation studies, we investigate the use of audio and score MIDI, the use of large-scale pretraining, the use of the cyclic training, and the use of the SiFiGAN post-processing.
@@ V-A Effectiveness of Serenade’s audio infilling task
We show that the proposed Serenade SiFiGAN framework had the highest overall N-MOS in the proposed systems. Despite the use of cyclic training to improve similarity scores, the N-MOS in Serenade was still very low due to the out-of-tune singing, possibly due to inaccuracies during MIDI estimation. However, Serenade SiFiGAN helped alleviate this by using the original F0 and correcting out-of-tune singing, resulting in a much higher naturalness score. We also observed that there was a slight degradation in naturalness in Serenade SiFiGAN compared to NU-SVC. This is a downside of the SiFiGAN postprocessing scheme, as simply replacing the F0 with the source’s would have likely introduced small mismatches in WORLD features.
@@ V-B Ablation studies
We first investigate compared the use of score MIDI labels provided in GTSinger (Ablation 1) with the use of the audio-extracted MIDI using the pretrained network [2023yongphonememidi] (Ablation 2). We found that using audio-extracted MIDI was better than using score MIDI, primarily because we opted to use audio-extracted MIDI during inference to keep the system practical for SSC. Nevertheless, we still found that using audio-extracted MIDI, and being more correlated to the input waveform, was more helpful in normalizing the pitch.
@@ V-C Correlations between objective and subjective tests
Finding a well-correlated objective metric can help future researchers validate ideas before running expensive subjective tests. With no previous works on SSC, we found that reference-based metric FAD [gui2024fad] using MERT [li2024mert] embeddings was the most correlated to the similarity scores, with a 0.93 correlation score. Nevertheless, large-scale evaluations with a multiple variety of systems still need to be conducted to find the best objective metric.
@@ VI Conclusions
We proposed Serenade, a novel framework to accomplish the SSC task and resolved the three main issues we found in this task, namely: singing style modeling, source style disentanglement, and melody retention. We showed a more generalized SSC framework and better modeling in breath features due to the audio infilling. Next, we used a cyclic training strategy to further disentangle the singing style of the source singer. Finally, we were able to alleviate out-of-tune singing by introducing a SiFiGAN vocoder post-processing method. We conclude that the Serenade framework showed better similarity scores in breathy and mixed singing styles than the baseline system but with a slight degradation in naturalness to do so.
@@ Acknowledgment
[…]